OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (anonymous):

\[610=x+y\]\[340=4x-2y\] I created a system of equations out of the information they gave me. From there, solve for one variable and plug into the other. Make sense?

OpenStudy (anonymous):

actually i dont really understand it at all

OpenStudy (anonymous):

Okay, so the first equation I wrote is the sum of all the one step and two step equations. They told us the sum of one step and two step equations is 610. For the second equation I again took the information they gave us. 4 times the amount of one step equations minus two times the number of two steps sums to 340. So x is the number of one step equations and y is number of two steps eqautions. That is where the two equations come from. To solve a two variable system of equations you need to isolate the variables in one of the equations and plug in for the other. So for this example: \[610-x=y\] plug this into the other equation 340=4x-2(610-x) Does that help?